# Integrable modules over quantum symmetric pair coideal subalgebras

### v2

@@ 3.3. Integrable modules
Let $\lambda,\mu\in X^{+}$. Then, as a $\mathbf{U}$-module, the $V^{\imath}(\lambda,\mu)$ is isomorphic to the quotient module of $\dot{\mathbf{U}}1_{w_{\bullet}\lambda+\mu}$ factored by the submodule generated by $\mathbf{R}^{+}1_{w_{\bullet}\lambda+\mu}$ and the set (3.3.1)
@@
Combining the two homomorphisms above, we see that $\mathbf{L}v^{\imath}_{\lambda,\mu}\simeq\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. By Lemma 3.3.2, the $\mathbf{U}$-module structure of $V^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}v^{\imath}_{\lambda,\mu}$. On the other hand, by [Wa23b, Proposition 3.4.3], the $\mathbf{U}$-module structure of $L^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. Hence, the assertion follows. ∎
@@
As explained in Section 1, the $\mathbf{U}$-modules $L^{\imath}(\lambda,\mu)$ in Proposition 3.3.3, which are isomorphic to $V^{\imath}(\lambda,\mu)$, are counterparts of ${}^{\omega}V(\lambda)\otimes V(\mu)$ in the theory of $\imath$canonical bases [BW18], [BW21], and the latter can be used to define integrable $\mathbf{U}$-modules. From this point of view, Definition 3.3.4 is quite natural.
@@ 4.1.1. Case $1:$ $\tau(k)=k=w_{\bullet}k$
First, let us consider the case where $\tau(k)=k=w_{\bullet}k$. For each $\zeta\in X^{\imath}$ and $\lambda\in X$ such that $\bar{\lambda}=\zeta$, the parity of the integer $\langle h_{k},\lambda\rangle$ is independent of the choice of $\lambda$. We call it the parity of $\zeta$ at $k$. Set $p_{k}(\zeta)\in\mathbb{Z}/2\mathbb{Z}=\{\bar{0},\bar{1}\}$ to be $\bar{0}$ if the parity of $\zeta$ at $k$ is even, and to be $\bar{1}$ otherwise. Similarly, given an integer $n$, set $p(n)$ to be $\bar{0}$ if $n$ is even, and to be $\bar{1}$ otherwise.
@@
Let us see where this definition comes from. Let $\mathbb{K}$ denote an algebraic closure of the field $\mathbb{Q}(q)$, and consider the following quadratic equations
@@
with variant $x$ and constants $a,b\in\mathbb{K}$. Since this equation is symmetric with respect to $x$ and $a$, it follows that if $c\in\mathbb{K}$ is a solution to $(\mathrm{QE})_{a,b}$, then $a$ is a solution to $(\mathrm{QE})_{c,b}$.
@@
Now, let us consider the $(n+1)$-dimensional irreducible representation $\mathbf{V}_{n}$ of $U_{q}(\mathfrak{sl}_{2})$, which we shall identify with the subalgebra of $\mathbf{U}$ generated by $E_{k}$, $F_{k}$, and $K_{k}^{\pm 1}$. By a similar argument to [Wa21, §3.1], one can show that the $B_{k}\in\mathbf{U}^{\imath}$ acts on $\mathbf{V}_{n}$ semisimply with eigenvalues $\kappa^{(n)},\kappa^{(n-2)},\dots,\kappa^{(-n)}$, where we set $\kappa^{(0)}:=\kappa_{k}$ and $b:=q_{k}\varsigma_{k}$. Therefore, we see that
@@ 4.1.2. Case $2:$ $\tau(k)=k\neq w_{\bullet}k$
Set $Y_{k}:=B_{k}-F_{k}$ and $Z_{k}:=F_{k}Y_{k}-q_{k}^{-2}Y_{k}F_{k}$. Then, the following hold (cf[BW21, §5.2]):
@@
Since $V$ is integrable, the $Y_{k}$, which is a scalar multiple of $T_{w_{\bullet}}(E_{\tau(k)})K_{k}^{-1}$, acts on $V$ locally nilpotently. In particular, there exists $y_{0}\geq 0$ such that
@@ 4.1.3. Case $3:$ $\tau(k)\neq k$
Set $Y_{k}:=B_{k}-F_{k}$. Then, the following hold (cf[BW21, §5.5]):
@@ 4.1.4. Summary
Let $V$ be a weight $\mathbf{U}$-module, $\lambda\in X$, $v\in V_{\lambda}$, and $k\in I_{\circ}$. Assume that $\langle h_{k},\lambda\rangle\geq 0$ and $E_{k}\mathbf{L}^{+}v=E_{\tau(k)}\mathbf{L}^{+}v=0$. Let $\mathbf{U}_{k}^{-}$ denote the subalgebra of $\mathbf{U}$ generated by $F_{k}$. Then, the following hold:
@@ 4.2. A Presentation of the $\mathbf{U}^{\imath}$-module $V^{\imath}(\lambda,\mu)$
Let $j\in I_{\bullet}$ and set $v:=F_{j}^{(\langle h_{j},\mu\rangle+1)}m^{\imath}_{\lambda,\mu}\in M^{\imath}(\lambda,\mu)$. Let $x\in\mathbf{L}^{+}$ and $k\in I_{\circ}$. By the triangular decomposition (2.4.1) of $\mathbf{L}$, the vector $E_{k}xv$ is, up to scalar multiplies, of the form $E_{k}x^{-}x^{+}m^{\imath}_{\lambda,\mu}$ with $x^{-}\in\mathbf{L}^{-}$, $x^{+}\in\mathbf{L}^{+}$. Since the $E_{k}$ commutes with the elements in $\mathbf{L}^{-}$, we have
@@
$\{\mathfrak{B}_{k,\overline{w_{\bullet}\lambda+\mu}+\mathrm{wt}^{\imath}(b)}^{(\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle+1)}bm^{\imath}_{\lambda,\mu}\mid k\in I_{\circ},\ b\in\mathbf{B}_{\mathbf{L}}(-\infty)\}.$
@@ 4.3. Some sufficient conditions
The assumption on $v$ implies that the $\mathbf{L}$-module $\mathbf{L}v$ is integrable. Since the $\mathbf{L}$ is a quantum group of finite type, every cyclic integrable module is finite-dimensional. Therefore, the set
@@ 5. Quantized coordinate algebras
  • the Cartan datum $I$ is of finite type.
  • the involution $\rho$ on $\mathbf{U}$ ([Lu93, 19.1.1]) preserves $\mathbf{U}^{\imath}$.
  • the canonical basis $\dot{\mathbf{B}}^{\imath}$ of $\dot{\mathbf{U}}^{\imath}$ possesses the stability property.
For example, the last two conditions above are satisfied if the parameters ${\boldsymbol{\varsigma}},{\boldsymbol{\kappa}}$ are as in [Wa23c, proof of Lemma 4.2.1] by [BW18, Proposition 4.6] and [Wa23c, Theorem 4.3.1].
@@ 5.1. Simple integrable modules
(cf[BW18, §6.6]). Since the $V^{\imath}(\lambda,\mu)$ is finite-dimensional, the existence of such a bilinear form ensures that it is semisimple. With some index set $\Gamma$, let
@@
Let $V$ be a simple integrable $\mathbf{U}^{\imath}$-module, and $v\in V$ a nonzero weight vector. Then, there exist $\lambda,\mu\in X^{+}$ and a $\mathbf{U}^{\imath}$-module homomorphism $V^{\imath}(\lambda,\mu)\rightarrow V$ which sends $v^{\imath}_{\lambda,\mu}$ to $v$. Since the $V$ is simple, this homomorphism must be surjective. The semisimplicity of $V^{\imath}(\lambda,\mu)$ implies that it has a simple submodule isomorphic to $V$. Therefore, the $V$ is isomorphic to $V(\gamma)$ for some $\gamma\in\Gamma$.
@@
Let $B(\gamma)=\{v^{\gamma}_{i}\mid i\in I_{\gamma}\}$ be a basis of $V(\gamma)$ consisting of weight vectors. Since the $V(\gamma)$ is finite-dimensional, the index set $I_{\gamma}$ is finite. By the integrability of $V(\gamma)$, for each $i\in I_{\gamma}$, we may take $\lambda_{i},\mu_{i}\in X^{+}$ such that there exists a $\mathbf{U}^{\imath}$-module homomorphism $V^{\imath}(\lambda_{i},\mu_{i})\rightarrow V(\gamma)$ which sends $v^{\imath}_{\lambda_{i},\mu_{i}}$ to $v^{\gamma}_{i}$.
@@ 5.2. Space of matrix coefficients
Given a linear space $V$, let $V^{*}$ denote the linear dual space $\operatorname{Hom}_{\mathbb{Q}(q)}(V,\mathbb{Q}(q))$ to $V$, and $\langle\mid\rangle\colon V\times V^{*}\rightarrow\mathbb{Q}(q)$ the canonical pairing.
@@
Let $V$ be an integrable cyclic $\mathbf{U}^{\imath}$-module with a cyclic vector $v$. Without any loss of generality, we may assume that $v$ is a weight vector. By the definition of integrable $\mathbf{U}^{\imath}$-modules, there exist $\lambda,\mu\in X^{+}$ and a $\mathbf{U}^{\imath}$-module homomorphism $V^{\imath}(\lambda,\mu)\rightarrow V$ which sends $v^{\imath}_{\lambda,\mu}$ to $v$. Since the $V$ is cyclic, the homomorphism must be surjective. Hence, the finite-dimensionality of $V^{\imath}(\lambda,\mu)$ implies that of $V$. Thus, we complete the proof. ∎
The following two statements can be proved by a similar way to classical results (e.g. see [P07, 6.2.6]).
@@
Let $b\in\dot{\mathbf{B}}^{\imath}$. Take $\lambda,\mu\in X^{+}$ in a way such that $bv^{\imath}_{\lambda,\mu}\neq 0$. Set $B:=\{b^{\prime}\in\dot{\mathbf{B}}^{\imath}\mid b^{\prime}v^{\imath}_{\lambda,\mu}\neq 0\}$. Note that the set $B$ is finite: Since the $\mathbf{U}^{\imath}$-module $V^{\imath}(\lambda,\mu)$ is integrable, there exist $(a_{j})_{j\in I_{\bullet}},(b_{j})_{j\in I_{\bullet}}\in\mathbb{Z}_{\geq 0}^{I_{\bullet}}$, $(c_{k})_{k\in I_{\circ}}\in\mathbb{Z}_{\geq 0}^{I_{\circ}}$ such that
@@
Next, we shall prove that the $\dot{\mathbf{B}}^{\imath*}$ spans $\mathbf{O}^{\imath}$. Let $\{v^{\gamma}_{i}\mid i\in I_{\gamma}\}$ be a linear basis of $V(\gamma)$ and $\{\phi^{\gamma}_{i}\mid i\in I_{\gamma}\}$ its dual basis. By the Peter-Weyl type decomposition 5.2.3, the $\mathbf{O}^{\imath}$ has the following linear basis:
@@
We can formally express this basis elements as
@@
The space of matrix coefficients for the semisimple integrable $\mathbf{U}^{\imath}$-modules coincides with the quantum coordinate algebra of $\dot{\mathbf{U}}^{\imath}$ in [BS22, 3.2.2].

### v3

@@ 3.3. Integrable modules
Let $\lambda,\mu\in X^{+}$. Then, the $\mathbf{U}$-module $V^{\imath}(\lambda,\mu)$ is isomorphic to the quotient module of $\dot{\mathbf{U}}1_{w_{\bullet}\lambda+\mu}$ factored by the submodule generated by $\mathbf{R}^{+}1_{w_{\bullet}\lambda+\mu}$ and the set (3.3.1)
@@
Combining the two homomorphisms above, we see that $\mathbf{L}v^{\imath}_{\lambda,\mu}\simeq\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. By Lemma 3.3.2, the $\mathbf{U}$-module structure of $V^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}v^{\imath}_{\lambda,\mu}$. On the other hand, by [25, Proposition 3.4.3], the $\mathbf{U}$-module structure of $L^{\imath}(\lambda,\mu)$ is determined by the $\mathbf{L}$-module structure of $\mathbf{L}(v_{w_{\bullet}\lambda}\otimes v_{\mu})$. Hence, the assertion follows. ∎
@@
As explained in Section 1, the $\mathbf{U}$-modules $L^{\imath}(\lambda,\mu)$ in Proposition 3.3.3, which are isomorphic to $V^{\imath}(\lambda,\mu)$, are counterparts of $V(\lambda,\mu)$ in the theory of $\imath$canonical bases [2], [3], and the latter can be used to define integrable $\mathbf{U}$-modules. From this point of view, Definition 3.3.4 is quite natural.
@@ 4.1.1. Case $1:$ $\tau(k)=k=w_{\bullet}k$
First, let us consider the case where $\tau(k)=k=w_{\bullet}k$. For each $\zeta\in X^{\imath}$ and $\lambda\in X$ such that $\bar{\lambda}=\zeta$, the parity of integer $\langle h_{k},\lambda\rangle$ is independent of the choice of $\lambda$. We call it the parity of $\zeta$ at $k$. Set $p_{k}(\zeta)\in\mathbb{Z}/2\mathbb{Z}=\{\bar{0},\bar{1}\}$ to be $\bar{0}$ if the parity of $\zeta$ at $k$ is even, and to be $\bar{1}$ otherwise. Similarly, for each integer $n$, set $p(n)$ to be $\bar{0}$ if $n$ is even, and to be $\bar{1}$ otherwise.
@@
Let us see where this definition comes from. Let $\mathbb{K}$ be an algebraic closure of the field $\mathbb{Q}(q)$, and consider the following quadratic equations
@@
with variable $x$ and constants $a,b\in\mathbb{K}$. Since this equation is symmetric with respect to $x$ and $a$, it follows that if $c\in\mathbb{K}$ is a solution to $(\mathrm{QE})_{a,b}$, then $a$ is a solution to $(\mathrm{QE})_{c,b}$.
@@
Now, let us consider the $(n+1)$-dimensional irreducible representation $\mathbf{V}_{n}$ of $U_{q}(\mathfrak{sl}_{2})$, which we identify with the subalgebra of $\mathbf{U}$ generated by $E_{k}$, $F_{k}$, and $K_{k}^{\pm 1}$. By a similar argument to [23, §3.1], one can show that the $B_{k}\in\mathbf{U}^{\imath}$ acts on $\mathbf{V}_{n}$ semisimply with eigenvalues $\kappa^{(n)},\kappa^{(n-2)},\dots,\kappa^{(-n)}$, where we set $\kappa^{(0)}:=\kappa_{k}$ and $b:=q_{k}\varsigma_{k}$. Therefore, we see that
@@ 4.1.2. Case $2:$ $\tau(k)=k\neq w_{\bullet}k$
Set $Y_{k}:=B_{k}-F_{k}$ and $Z_{k}:=F_{k}Y_{k}-q_{k}^{-2}Y_{k}F_{k}$. Then, the following hold (cf[3, §5.2]):
@@
Since $V$ is integrable, the element $Y_{k}$, which is a scalar multiple of $T_{w_{\bullet}}(E_{\tau(k)})K_{k}^{-1}$, acts on $V$ locally nilpotently. In particular, there exists $y_{0}\geq 0$ such that
@@ 4.1.3. Case $3:$ $\tau(k)\neq k$
Set $Y_{k}:=B_{k}-F_{k}$. Then, the following hold (cf[3, §5.5]):
@@ 4.1.4. Summary
Let $V$ be a weight $\mathbf{U}$-module, $\lambda\in X$, $v\in V_{\lambda}$, and $k\in I_{\circ}$. Assume that $\langle h_{k},\lambda\rangle\geq 0$ and $E_{k}\mathbf{L}^{+}v=E_{\tau(k)}\mathbf{L}^{+}v=0$. Let $\mathbf{U}_{k}^{-}$ denote the subalgebra of $\mathbf{U}$ generated by $F_{k}$. Then, the following hold.
@@ 4.2. A Presentation of the $\mathbf{U}^{\imath}$-module $V^{\imath}(\lambda,\mu)$
Let $j\in I_{\bullet}$ and set $v:=F_{j}^{(\langle h_{j},\mu\rangle+1)}m^{\imath}_{\lambda,\mu}\in M^{\imath}(\lambda,\mu)$. Let $x\in\mathbf{L}^{+}$ and $k\in I_{\circ}$. By the triangular decomposition (2.4.1) of $\mathbf{L}$, the vector $E_{k}xv$ is, up to scalar multiplies, of the form $E_{k}x^{-}x^{+}m^{\imath}_{\lambda,\mu}$ with $x^{-}\in\mathbf{L}^{-}$, $x^{+}\in\mathbf{L}^{+}$. Since $E_{k}$ commutes with the elements in $\mathbf{L}^{-}$, we have
@@
$\{\mathfrak{B}_{k,\overline{w_{\bullet}\lambda+\mu+\mathrm{wt}(b)}}^{(\langle h_{k},w_{\bullet}\lambda+\mu+\operatorname{wt}(b)\rangle+1)}bm^{\imath}_{\lambda,\mu}\mid k\in I_{\circ},\ b\in\mathbf{B}_{\mathbf{L}}(-\infty)\}.$
@@ 4.3. Some sufficient conditions
The assumption on $v$ implies that the $\mathbf{L}$-module $\mathbf{L}v$ is integrable. Since $\mathbf{L}$ is a quantum group of finite type, every cyclic integrable module is finite-dimensional. Therefore, the set
@@ 5. Quantized coordinate algebras
  • the Cartan datum $I$ is of finite type,
  • the involution $\rho$ on $\mathbf{U}$ ([16, 19.1.1]) given by
    $\rho(E_{i})=q_{i}K_{i}F_{i},\quad\rho(F_{i})=q_{i}K_{i}^{-1}E_{i},\quad\rho(K_{h})=K_{h},$
    preserves $\mathbf{U}^{\imath}$,
  • the canonical basis $\dot{\mathbf{B}}^{\imath}$ of $\dot{\mathbf{U}}^{\imath}$ possesses the stability property (or strong compatibility) (cf[2, Remark 6.18]).
For example, if $\kappa_{i}$ are all zero and each $\varsigma_{i}$ is a certain signed power of $q_{i}$ as in [26, proof of Lemma 4.2.1], then the last two conditions above are satisfied by [2, Proposition 4.6] and [26, Theorem 4.3.1]. The parameters in [26, proof of Lemma 4.2.1] are almost the same as in [2, Table 3], where the constants are deduced by requiring that each $\varsigma_{i}$ is a signed power of $q_{i}$ and that the bar-involution (which was used to construct $\dot{\mathbf{B}}^{\imath}$) on $\mathbf{U}^{\imath}$ exists.
@@ 5.1. Simple integrable modules
(cf[2, §6.6]). Since $V^{\imath}(\lambda,\mu)$ is finite-dimensional, the existence of such a bilinear form ensures that it is semisimple. In fact, if $W$ is a submodule of $V^{\imath}(\lambda,\mu)$, and $W^{\perp}$ denotes the orthogonal complement of $W$ with respect to $(,)_{\lambda,\mu}$, then $W^{\perp}$ is also a submodule of $V^{\imath}(\lambda,\mu)$ since $\rho$ preserves $\mathbf{U}^{\imath}$. With some index set $\Gamma$, let
@@
Let $V$ be a simple integrable $\mathbf{U}^{\imath}$-module, and $v\in V$ a nonzero weight vector. Then, there exist $\lambda,\mu\in X^{+}$ and a $\mathbf{U}^{\imath}$-module homomorphism $V^{\imath}(\lambda,\mu)\rightarrow V$ which sends $v^{\imath}_{\lambda,\mu}$ to $v$. Since $V$ is simple, this homomorphism must be surjective. The semisimplicity of $V^{\imath}(\lambda,\mu)$ implies that it has a simple submodule isomorphic to $V$. Therefore, $V$ is isomorphic to $V(\gamma)$ for some $\gamma\in\Gamma$.
@@
Let $B(\gamma)=\{v^{\gamma}_{i}\mid i\in I_{\gamma}\}$ be a basis of $V(\gamma)$ consisting of weight vectors. Since $V(\gamma)$ is finite-dimensional, the index set $I_{\gamma}$ is finite. By the integrability of $V(\gamma)$, for each $i\in I_{\gamma}$, we may take $\lambda_{i},\mu_{i}\in X^{+}$ such that there exists a $\mathbf{U}^{\imath}$-module homomorphism $V^{\imath}(\lambda_{i},\mu_{i})\rightarrow V(\gamma)$ which sends $v^{\imath}_{\lambda_{i},\mu_{i}}$ to $v^{\gamma}_{i}$.
@@ 5.2. Space of matrix coefficients
Given a linear space $V$, let $V^{*}$ denote the linear dual space $\operatorname{Hom}_{\mathbb{Q}(q)}(V,\mathbb{Q}(q))$, and $\langle\ \mid\ \rangle\colon V\times V^{*}\rightarrow\mathbb{Q}(q)$ the canonical pairing.
@@
Let $V$ be an integrable cyclic $\mathbf{U}^{\imath}$-module with a cyclic vector $v$. Without any loss of generality, we may assume that $v$ is a weight vector. By the definition of integrable $\mathbf{U}^{\imath}$-modules, there exist $\lambda,\mu\in X^{+}$ and a $\mathbf{U}^{\imath}$-module homomorphism $V^{\imath}(\lambda,\mu)\rightarrow V$ which sends $v^{\imath}_{\lambda,\mu}$ to $v$. Since $V$ is cyclic, the homomorphism must be surjective. Hence, the finite-dimensionality of $V^{\imath}(\lambda,\mu)$ implies that of $V$. Thus, we complete the proof. ∎
The following two statements can be proved by a similar way to classical results (e.g. see [20, 6.2.6]).
@@
Let $b\in\dot{\mathbf{B}}^{\imath}$. Take $\lambda,\mu\in X^{+}$ in a way such that $bv^{\imath}_{\lambda,\mu}\neq 0$. Set
$B:=\{b^{\prime}\in\dot{\mathbf{B}}^{\imath}\mid b^{\prime}v^{\imath}_{\lambda,\mu}\neq 0\}.$
Note that the set $B$ is finite. In fact, since the $\mathbf{U}^{\imath}$-module $V^{\imath}(\lambda,\mu)$ is integrable, there exist $(a_{j})_{j\in I_{\bullet}},(b_{j})_{j\in I_{\bullet}}\in\mathbb{Z}_{\geq 0}^{I_{\bullet}}$, $(c_{k})_{k\in I_{\circ}}\in\mathbb{Z}_{\geq 0}^{I_{\circ}}$ such that
@@
Next, we shall prove that the $\dot{\mathbf{B}}^{\imath*}$ spans $\mathbf{O}^{\imath}$. Let $\{v^{\gamma}_{i}\mid i\in I_{\gamma}\}$ be a linear basis of $V(\gamma)$ and $\{\phi^{\gamma}_{i}\mid i\in I_{\gamma}\}$ its dual basis. By the Peter-Weyl type decomposition 5.2.3, $\mathbf{O}^{\imath}$ has the following linear basis:
@@
We can formally express these basis elements as
@@
The space of matrix coefficients for the semisimple integrable $\mathbf{U}^{\imath}$-modules coincides with the quantum coordinate algebra of $\dot{\mathbf{U}}^{\imath}$ in [1, 3.2.2].